# Hierarchical Empowerment: Towards Tractable Empowerment-Based Skill-Learning

### v1

@@ Abstract
General purpose agents will require large repertoires of skills. Empowerment—the maximum mutual information between skills and the states—provides a pathway for learning large collections of distinct skills, but mutual information is difficult to optimize. We introduce a new framework, Hierarchical Empowerment, that makes computing empowerment more tractable by integrating concepts from Goal-Conditioned Hierarchical Reinforcement Learning. Our framework makes two specific contributions. First, we introduce a new variational lower bound on mutual information that can be used to compute empowerment over short horizons. Second, we introduce a hierarchical architecture for computing empowerment over exponentially longer time scales. We verify the contributions of the framework in a series of simulated robotics tasks. In a popular ant navigation domain, our four level agents are able to learn skills that cover a surface area over two orders of magnitude larger than prior work.
@@ 1 Introduction
General purpose robots that can execute a wide array of skills have the potential to greatly improve human productivity. Yet the question of how to enable agents to learn large sets of skills remains a central challenge in artificial intelligence research. Empowerment [Klyubin et al., 2005, Salge et al., 2014, Jung et al., 2012, Gregor et al., 2017] provides a compelling pathway for implementing such general purpose robots. As the mutual information between skills and the states to which they lead, empowerment enables agents to learn the largest space of skills such that each skill can reliably reach a distinct area of the state space. A major problem with using empowerment for skill-learning is that it is difficult to compute. Recent work [Mohamed and Rezende, 2015, Gregor et al., 2017, Eysenbach et al., 2019] has made some progress using Reinforcement Learning (RL) [Sutton and Barto, 1998] to optimize a variational lower bound on empowerment, but these methods are limited to learning small spaces of skills.
There are two problems with the current approach of using RL to compute empowerment. The first issue is the reward function employed by existing methods does not incentivize skills to target particular regions of the state space when there is significant overlap among the skills. For instance, at the start of training, when most of the skill space generally visits the same states near the starting state, the reward function used by many empowerment-based skill-learning methods will not encourage the skills to differentiate. The second problem is that it will be difficult to use RL to compute long-term empowerment as RL has consistently struggled at learning long horizon tasks [Andrychowicz et al., 2017, Nair et al., 2018, Trott et al., 2019, Levy et al., 2019, Nachum et al., 2018].
Central to our work is the insight from Choi et al. [2021] that goal-conditioned RL can be formulated as a variational lower bound on the mutual information between skills and states. Goal-Conditioned RL offers a promising lower bound on mutual information because it uses a reward function that explicitly encourages skills to specialize and target particular states. However, Choi et al. [2021] do not provide a way to learn the distribution of goal states (i.e., skills) used in goal-conditioned RL and instead require a hand-crafted space of goal states. Further, the authors do not offer a practical approach for computing long-term empowerment.
@@
We introduce a framework, Hierarchical Empowerment, that takes a step toward tractable long-term empowerment computation. The framework makes two contributions. The first contribution, Goal-Conditioned Empowerment, is a new variational lower bound objective on empowerment that extends Choi et al. [2021] by learning the space of achievable goal states. We show that after applying the reparameterization trick [Kingma and Welling, 2014], our mutual information objective with respect to the learned goal space is a maximum entropy [Ziebart et al., 2008] bandit problem, in which the goal space is rewarded for being large and containing achievable goals. Optimizing the full Goal-Conditioned Empowerment objective takes the form of an automated curriculum of goal-conditioned RL problems. Figure 1(Left) illustrates this idea.
@@
We evaluate the two proposed contributions of our framework in a series of simulated robotic navigation tasks. The experimental results support both contributions. Goal-Conditioned Empowerment outperforms the baseline at learning the short-term empowerment of states. In addition, our results indicate that hierarchy is needed to compute long-term empowerment as agents with more levels of hierarchy outperformed those with fewer. A video presentation of our results is available at the following url: https://www.youtube.com/watch?v=kwLUYMsMRNI.
@@ 2.1 Goal-Conditioned Markov Decision Processes (MDPs)
Central to our work is the concept of Goal-Conditioned MDPs [Kaelbling, 1993, Schaul et al., 2015], which represent a decision making setting in which an agent needs to learn a distribution of tasks. We define Goal-Conditioned MDPs by the tuple $\{\mathcal{S},p(s_{0}),\mathcal{G},p(g),\mathcal{A},T(s_{t+1}|s_{t},a_{t}),\epsilon,r(s_{t},g,a_{t})\}$. $\mathcal{S}$ is the space of states. $p(s_{0})$ is the distribution of starting states. $\mathcal{G}$ is the space of goals and is assumed to be a subset of $\mathcal{S}$. $p(g)$ is the distribution of goals for the agent to achieve. $\mathcal{A}$ is the action space. $T(s_{t+1}|s_{t},a_{t})$ is the transition dynamics of the environment. We assume episodes terminate in an absorbing state when the agent is within an $\epsilon$-ball of the goal state. That is, for some distance function $d(\cdot)$, if $d(s_{t},g)<\varepsilon$, then $p(s_{t+1}|s_{t},a_{t},g)=\delta$, in which $\delta$ is the impulse function, when $s_{t+1}=s_{t}$. Otherwise, if $d(s_{t},g)\geq\epsilon$, $p(s_{t+1}|s_{t},a_{t},g)=T(s_{t+1}|s_{t},a_{t})$. $r(s_{t},g,a_{t})$ is the reward function for a (state,goal,action) tuple. A policy $\pi_{\theta}$ in a Goal-Conditioned MDP is a mapping from states and goals to a distribution over the action space.
The objective in a goal-conditioned MDP is learn the parameters $\theta$ of the policy that maximize the sum of discounted rewards averaged over all goals and goal-conditioned trajectories:
$\max_{\theta}\mathrm{E}_{g\sim p(g),\tau\sim p(\tau|g)}[\sum_{t=0}^{\infty}\gamma^{t}r(s_{t},g,a_{t})],$
in which $\tau$ is a trajectory of states and actions $(s_{0},a_{0},s_{1},a_{1},\dots)$.
@@ 2.2 The Skill Channel and Empowerment
Key to our approach for skill-learning is the observation that skills are an instance of noisy channels from Information Theory [Cover and Thomas, 2006]. We will refer to this noisy channel as the skill channel. The skill channel is defined by the tuple $\{\mathcal{Z},p(s_{n}|s_{0},z),\mathcal{S}\}$, in which $\mathcal{Z}$ is the set of skills that can be executed from a start state $s_{0}$, $\mathcal{S}$ is the set of states, $n$ is the number of actions contained within a skill, and $p(s_{n}|s_{0},z)$ is the probability of a skill $z$ that started in state $s_{0}$ terminating in state $s_{n}$. The distribution $p(s_{n}|s_{0},z)$ depends on the skill-conditioned policy $\pi(a_{t}|s_{t},z)$ and the transition dynamics of the environment.
The connection between skills and noisy channels is important because of the discovery by Shannon [1948] that the largest number of distinct inputs (in bits) that can be sent over a channel (also referred to as the capacity of the channel) is the maximum mutual information of the channel. In this context, distinct means the input to the channel can be identified by the output with arbitrarily small probability of error. Mutual information is thus an appealing objective for implementing general purpose agents that need to reliably reach many specific states because maximizing the mutual information of the skill channel will find the largest set of skills such that each skill targets a unique area of the state space.
The channel capacity of the skill channel is known as Empowerment [Klyubin et al., 2005], which is thus defined as
$\mathcal{E}(s_{0})$ $=\max_{p(z),\pi}I(Z;S_{n}|s_{0}).$ (1)
The mutual information of the skill channel is accordingly defined as
$I(Z;S_{n}|s_{0})=H(Z|s_{0})-H(Z|s_{0},S_{n}).$ (2)
Note that in the literature empowerment is often defined with respect to more simple skills, such as single primitive actions or open loop action sequences [Klyubin et al., 2005, Mohamed and Rezende, 2015], but we use the more general definition developed by Gregor et al. [2017] that permits learnable closed loop skills.
@@ 2.3 Challenges with Computing Empowerment
One major difficulty with calculating empowerment is that computing the posterior distribution $p(z|s_{0},s_{n})$, located within the conditional entropy term $H(Z|s_{0},S_{n})$, is often intractable. Analytically calculating the posterior distribution would require integrating over the intermediate actions and states $a_{1},s_{1},a_{2}\dots,s_{n-1},a_{n-1}$. Inspired by Mohamed and Rezende [2015], many empowerment-based skill-learning approaches [Gregor et al., 2017, Eysenbach et al., 2019, Zhang et al., 2021, Strouse et al., 2021, Warde-Farley et al., 2018, Baumli et al., 2020, Achiam et al., 2018, Hansen et al., 2020] attempt to overcome the issue by optimizing a variational lower bound on the mutual information of the skill channel. In this lower bound, a learned variational distribution $q_{\psi}(z|s_{0},s_{n})$ parameterized by $\psi$ is used in place of the problematic posterior $p(z|s_{0},s_{n})$. (See Barber and Agakov [2003] for original proof of the lower bound.) The variational lower bound is accordingly defined
$I^{VB}(Z;S_{n}|s_{0})$ $=H(Z|s_{0})+\mathrm{E}_{z\sim p(z|s_{0}),s_{n}\sim p_{\theta}(s_{n}|s_{0},z)}[\log q_{\psi}(z|s_{0},s_{n})],$ (3)
@@
The common approach in empowerment-based skill-learning is to train the learned variational distribution $q_{\psi}$ with maximum likelihood learning so that the variational distribution approaches the true posterior. The skill distribution $p(z)$ is typically held fixed so $H(Z|s_{0})$ is constant and can be ignored. The last term in the variational lower bound, as observed by [Gregor et al., 2017], is a Reinforcement Learning problem in which the reward for a (state,skill,action) tuple is $r(s_{t},z,a_{t})=0$ for the first $n-1$ actions and then $r(s_{n-1},z,a_{n-1})=\mathrm{E}_{s_{n}\sim p(s_{n}|s_{n-1},a_{n})}[\log q_{\psi}(z|s_{0},s_{n}))]$ for the final action.
Unfortunately, this reward function does not encourage skills to differentiate and target distinct regions of the state space (i.e., increase mutual information) when the skills are currently overlapping. For instance, at the start of training when skills tend to visit the same states in the immediate vicinity of the start state, the reward $q_{\psi}(z|s_{0},s_{n})$ will typically be small for most $(s_{0},z,s_{n})$ tuples. The undifferentiated reward will in turn produce little change in the skill-conditioned policy, which then triggers little change in the learned variational distribution and reward $q_{\psi}$, making it difficult to maximize mutual information.
@@ 2.4 Goal-Conditioned RL as a Mutual Information Variational Lower Bound
Choi et al. [2021] observed that the goal-conditioned RL objective below is also a variational lower bound on mutual information.
$\mathrm{E}_{g\sim p(g|s_{0})}[\mathrm{E}_{s_{1},\dots,s_{n}\sim p_{\theta}(s_{1},\dots,s_{n}|s_{0},g)}[\log q(g|s_{0},s_{n})]],$ (4)
@@
The expression in 4 is a goal-conditioned RL objective as there is an outer expectation over a fixed distribution of goal states $p(g|s_{0})$ and an inner expectation with respect to the distribution of trajectories produced by a goal state. The reward function is 0 for all time steps except for the last when it is the $\log$ probability of the skill $g$ sampled from a Gaussian distribution centered on the skill-terminating state $s_{n}$ with fixed standard deviation $\sigma_{0}$ selected by the designer. Like any goal-conditioned RL reward function, this reward encourages goal-conditioned policy $\pi_{\theta}(s_{t},g)$ to target the conditioned goal as the reward is higher when the goal $g$ is near the skill-terminating state $s_{n}$.
The goal-conditioned objective in Equation 4, which reduces to $\mathrm{E}_{g\sim p(g|s_{0}),s_{n}\sim p(s_{n}|s_{0},g)}[\log q(g|s_{0},s_{n})]$, is also a variational lower bound on the mutual information of the skill channel. The distribution of skills is represented by a distribution of goal states, and a variational distribution—in this case, a fixed variance Gaussian distribution—is used in place of the true posterior $p(g|s_{0},s_{n})$. This variational lower bound is appealing because it explicitly encourages skills to target specific states. Even when there is significant overlap among the skills in regards to the states visited, goal-conditioned RL will still encourage skills to specialize and achieve distinct states.
One issue with using goal-conditioned RL as a variational lower bound on mutual information is the hand-crafted distribution of skills $p(g|s_{0})$. If this is significantly smaller or larger than the region of states that can be reached in $n$ actions by the goal-conditioned policy, then maximizing the goal-conditioned RL objective would only serve as a loose lower bound on empowerment because the objective would either learn fewer skills than it is capable of or too many redundant skills. Another issue with the objective is that it is not a practical objective for learning lengthy goal-conditioned policies given that RL has consistently struggled at performing credit assignment over long horizons using temporal difference learning [Levy et al., 2019, Nachum et al., 2018, McClinton and Levy, 2021]. Thus, goal-conditioned RL is not an appealing objective for computing long-term empowerment.

### v2

# Hierarchical Empowerment: Towards Tractable Empowerment-Based Skill Learning

@@ Abstract
General purpose agents will require large repertoires of skills. Empowerment—the maximum mutual information between skills and states—provides a pathway for learning large collections of distinct skills, but mutual information is difficult to optimize. We introduce a new framework, Hierarchical Empowerment, that makes computing empowerment more tractable by integrating concepts from Goal-Conditioned Hierarchical Reinforcement Learning. Our framework makes two specific contributions. First, we introduce a new variational lower bound on mutual information that can be used to compute empowerment over short horizons. Second, we introduce a hierarchical architecture for computing empowerment over exponentially longer time scales. We verify the contributions of the framework in a series of simulated robotics tasks. In a popular ant navigation domain, our four level agents are able to learn skills that cover a surface area over two orders of magnitude larger than prior work.
@@ 1 Introduction
How to implement general purpose agents that can execute a vast array of skills remains a central challenge in artificial intelligence research. Empowerment (Klyubin et al., 2005; Salge et al., 2014; Jung et al., 2012; Gregor et al., 2017) offers a compelling objective for equipping agents with a large set of skills. As the maximum mutual information (MI) between skills and the states to which they lead, empowerment enables agents to learn large sets of distinct skills that each target a different area of the state space. The major problem with using empowerment for skill learning is that mutual information is difficult to optimize.
Recent work (Mohamed & Rezende, 2015; Gregor et al., 2017; Eysenbach et al., 2019) has made some progress using Reinforcement Learning (RL) (Sutton & Barto, 1998) to optimize a variational lower bound on mutual information, but these methods are limited to learning small spaces of skills. The main issue is that the reward function produced by the conventional MI variational lower bound does not incentivize skills to target particular regions of the state space when there is significant overlap among the skills. A potentially better alternative is to use Goal-Conditioned Reinforcement Learning (GCRL) as the variational lower bound to MI (Choi et al., 2021). The GCRL variational lower bound employs a reward function that always encourages skills to target specific states regardless of the degree to which the skills are overlapping. However, a key problem with GCRL is that it requires a hand-crated goal space. Not only does this require domain expertise, but a hand-crafted goal space may only achieve a loose lower bound on empowerment if the goal space is too small or too large as the agent may learn too few skills or too many redundant skills. In addition, it is unclear whether goal-conditioned RL can be used to compute long-term empowerment (i.e., learn long-horizon skills) given the poor performance of RL in long horizon settings (Andrychowicz et al., 2017; Nair et al., 2018; Trott et al., 2019; Levy et al., 2019; Nachum et al., 2018).
@@
We introduce a framework, Hierarchical Empowerment, that takes a step toward tractable long-term empowerment computation. The framework makes two contributions. The first contribution, Goal-Conditioned Empowerment, is a new variational lower bound objective on empowerment that improves on GCRL by learning the space of achievable goal states. We show that after applying the reparameterization trick (Kingma & Welling, 2014), our mutual information objective with respect to the learned goal space is a maximum entropy (Ziebart et al., 2008) bandit problem, in which the goal space is rewarded for being large and containing achievable goals. Optimizing the full Goal-Conditioned Empowerment objective takes the form of an automated curriculum of goal-conditioned RL problems. Figure 1(Left) illustrates this idea.
@@
Note that our framework has two significant limitations. First, similar to other empowerment-based skill-learning methods, our framework assumes the agent has access to a model of the environment’s transition dynamics. Second, the framework assumes the learned goal space is a uniform distribution over a subset of the state space, which significantly limits the types of domains to which the framework can be applied. We discuss these limitations in more detail in section 3.
We evaluate the two proposed contributions of our framework in a series of simulated robotic navigation tasks. The experimental results support both contributions. Goal-Conditioned Empowerment outperforms the baselines at learning the short-term empowerment of states. In addition, our results indicate that hierarchy is useful for computing long-term empowerment as agents with more levels of hierarchy outperformed those with fewer. In our largest domain, our agents are able to complete navigation tasks that are orders of magnitude longer than related work. A video presentation of our results is available at the following url: https://www.youtube.com/watch?v=OOjtC30VjPQ.
@@ 2.1 Goal-Conditioned Markov Decision Processes (MDPs)
A core concept in our work is Goal-Conditioned MDPs (Kaelbling, 1993; Schaul et al., 2015), which represent a decision making setting in which an agent needs to learn a distribution of tasks. We define Goal-Conditioned MDPs by the tuple $\{\mathcal{S},p(s_{0}),\mathcal{G},p(g),\mathcal{A},T(s_{t+1}|s_{t},a_{t}),r(s_{t},g,a_{t})\}$. $\mathcal{S}$ is the space of states; $p(s_{0})$ is the distribution of starting states; $\mathcal{G}$ is the space of goals and is assumed to be a subset of $\mathcal{S}$; $p(g)$ is the distribution of goals for the agent to achieve; $\mathcal{A}$ is the action space; $T(s_{t+1}|s_{t},a_{t})$ is the transition dynamics of the environment; $r(s_{t},g,a_{t})$ is the reward function for a (state,goal,action) tuple. A policy $\pi_{\theta}(\cdot|s_{t},z)$ in a Goal-Conditioned MDP is a mapping from states and goals to a distribution over the action space.
The objective in a goal-conditioned MDP is to learn the parameters $\theta$ of the policy that maximize the sum of discounted rewards averaged over all goals and goal-conditioned trajectories:
$\operatorname*{arg\,max}_{\theta}\mathbb{E}_{g\sim p(g),\tau\sim p(\tau|g)}[\sum_{t=0}^{\infty}\gamma^{t}r(s_{t},g,a_{t})],$
in which $\tau$ is a trajectory of states and actions $(s_{0},a_{0},s_{1},a_{1},\dots)$. The goal-conditioned MDP objective can be optimized with Reinforcement Learning, such as policy gradient methods (Sutton et al., 1999a).
@@ 2.2 The Skill Channel and Empowerment
Central to our approach for skill learning is the observation that skills are an instance of noisy channels from Information Theory (Cover & Thomas, 2006). The proposed skill is the message input into the noisy channel. The state that results from executing the skill is the output of the channel. We will refer to this noisy channel as the skill channel. The skill channel is defined by the tuple $\{\mathcal{S},\mathcal{Z},p(s_{n}|s_{0},z)\}$, in which $\mathcal{S}$ is the set of states, $\mathcal{Z}$ is the set of skills that can be executed from a start state $s_{0}\in\mathcal{S}$, $n$ is the number of actions contained within a skill, and $p(s_{n}|s_{0},z)$ is the probability of a skill $z$ that started in state $s_{0}$ terminating in state $s_{n}$ after $n$ actions. That is, $p(s_{n}|s_{0},z)=\int_{a_{0},s_{1},a_{1},\dots,s_{n-1},a_{n-1}}\pi(a_{0}|s_{0},z)p(s_{1}|s_{0},a_{0})\pi(a_{1}|s_{1},z)\dots\pi(a_{n-1}|s_{n-1},z)p(s_{n}|s_{n-1},a_{n-1})$. The distribution $p(s_{n}|s_{0},z)$ thus depends on the skill-conditioned policy $\pi(a_{t}|s_{t},z)$ and the transition dynamics of the environment $p(s_{t+1}|s_{t},a_{t})$.
The mutual information of the skill channel, $I(Z;S_{n}|s_{0})$, describes the number of distinct skills that can be executed from state $s_{0}$ given the current skill distribution $p(z|s_{0})$ and the skill-conditioned policy $\pi(a_{t}|s_{t},z)$. Mathematically, the mutual information of the skill channel can be defined:
$I(Z;S_{n}|s_{0})=H(Z|s_{0})-H(Z|s_{0},S_{n}).$
That is, the number of distinct skills that can be executed from state $s_{0}$ will be larger when (i) the entropy $H(Z|s_{0})$ (i.e., the size) of the skill space is larger and/or (ii) the conditional entropy $H(Z|s_{0},S_{n})$ is smaller (i.e., the skills are more differentiated as they target distinct regions of the state space).
Given our goal of equipping agents with as large a skillset as possible, the purpose of this paper is to develop an algorithm that can learn a distribution over skills $p_{\phi}(z|s_{0})$ and a skill-conditioned policy $\pi_{\theta}(s_{t},z)$ that maximize the mutual information of the skill channel:
$\operatorname*{arg\,max}_{p_{\phi},\pi_{\theta}}I(Z;S_{n}|s_{0})$ (1)
The maximum mutual information (equivalently, the channel capacity) of the skill channel is known as Empowerment (Klyubin et al., 2005). Note that in the literature empowerment is often defined with respect to more simple skills, such as single primitive actions or open loop action sequences (Klyubin et al., 2005; Mohamed & Rezende, 2015), but we use the more general definition developed by Gregor et al. (2017) that permits learnable closed loop skills.
@@ 2.3 Challenges with Computing Empowerment
One major difficulty with calculating empowerment (i.e., maximizing the mutual information of the skill channel) is that computing the posterior distribution $p(z|s_{0},s_{n})$, located within the conditional entropy term $H(Z|s_{0},S_{n})$, is often intractable. Analytically calculating the posterior distribution would require integrating over the intermediate actions and states $a_{0},s_{1},a_{1}\dots,s_{n-1},a_{n-1}$. Inspired by Mohamed & Rezende (2015), many empowerment-based skill-learning approaches (Gregor et al., 2017; Eysenbach et al., 2019; Zhang et al., 2021; Strouse et al., 2021; Warde-Farley et al., 2018; Baumli et al., 2020; Achiam et al., 2018; Hansen et al., 2020) attempt to overcome the issue by optimizing a variational lower bound on the mutual information of the skill channel. In this lower bound, a learned variational distribution $q_{\psi}(z|s_{0},s_{n})$ parameterized by $\psi$ is used in place of the problematic posterior $p(z|s_{0},s_{n})$. (See Barber & Agakov (2003) for original proof of the lower bound.) The variational lower bound is accordingly defined
$I^{VB}(Z;S_{n}|s_{0})$ $=H(Z|s_{0})+\mathbb{E}_{z\sim p(z|s_{0}),s_{n}\sim p_{\theta}(s_{n}|s_{0},z)}[\log q_{\psi}(z|s_{0},s_{n})],$ (2)
@@
The common approach in empowerment-based skill learning is to train the learned variational distribution $q_{\psi}$ with maximum likelihood learning so that the variational distribution approaches the true posterior. The skill distribution $p(z)$ is typically held fixed so $H(Z|s_{0})$ is constant and can be ignored. The last term in the variational lower bound, as observed by (Gregor et al., 2017), is a Reinforcement Learning problem in which the reward for a (state,skill,action) tuple is $r(s_{t},z,a_{t})=0$ for the first $n-1$ actions and then $r(s_{n-1},z,a_{n-1})=\mathbb{E}_{s_{n}\sim p(s_{n}|s_{n-1},a_{n})}[\log q_{\psi}(z|s_{0},s_{n}))]$ for the final action.
Unfortunately, this reward function does not encourage skills to differentiate and target distinct regions of the state space (i.e., increase mutual information) when the skills overlap and visit similar states. For instance, at the start of training when skills tend to visit the same states in the immediate vicinity of the start state, the reward $q_{\psi}(z|s_{0},s_{n})$ will typically be small for most $(s_{0},z,s_{n})$ tuples. The undifferentiated reward will in turn produce little change in the skill-conditioned policy, which then triggers little change in the learned variational distribution and reward $q_{\psi}$, making it difficult to maximize mutual information.
@@ 2.4 Goal-Conditioned RL as a Mutual Information Variational Lower Bound
Choi et al. (2021) observed that goal-conditioned RL, in the form of the objective below, is also a variational lower bound on mutual information.
$\mathbb{E}_{g\sim p(g|s_{0})}[\mathbb{E}_{s_{1},\dots,s_{n}\sim p_{\theta}(s_{1},\dots,s_{n}|s_{0},g)}[\log q(g|s_{0},s_{n})]],$ (3)
@@
That is, $q(\cdot|s_{0},s_{n})$ is a diagonal Gaussian distribution with mean $s_{n}$ and a fixed standard deviation $s_{0}$ set by the designer. The expression in 3 is a goal-conditioned RL objective as there is an outer expectation over a fixed distribution of goal states $p(g|s_{0})$ and an inner expectation with respect to the distribution of trajectories produced by a goal state. The reward function is 0 for all time steps except for the last when it is the $\log$ probability of the skill $g$ sampled from a Gaussian distribution centered on the skill-terminating state $s_{n}$ with fixed standard deviation $\sigma_{0}$. Like any goal-conditioned RL reward function, this reward encourages goal-conditioned policy $\pi_{\theta}(s_{t},g)$ to target the conditioned goal as the reward is higher when the goal $g$ is near the skill-terminating state $s_{n}$.
The goal-conditioned objective in Equation 3, which reduces to $\mathbb{E}_{g\sim p(g|s_{0}),s_{n}\sim p(s_{n}|s_{0},g)}[\log q(g|s_{0},s_{n})]$, is also a variational lower bound on the mutual information of the skill channel. In this case, the variational distribution is a fixed variance Gaussian distribution. Unlike the typical variational lower bound in empowerment-based skill learning, goal-conditioned RL uses a reward function that explicitly encourages skills to target specific states, even when there is significant overlap among the skills in regards to the states visited.
One issue with using goal-conditioned RL as a variational lower bound on mutual information is the hand-crafted distribution of skills $p(g|s_{0})$. In addition to requiring domain expertise, if this is significantly smaller or larger than the region of states that can be reached in $n$ actions by the goal-conditioned policy, then maximizing the goal-conditioned RL objective would only serve as a loose lower bound on empowerment because the objective would either learn fewer skills than it is capable of or too many redundant skills. Another issue with the objective is that it is not a practical objective for learning lengthy goal-conditioned policies given that RL has consistently struggled at performing credit assignment over long horizons using temporal difference learning (Levy et al., 2019; Nachum et al., 2018; McClinton et al., 2021).